OpenStudy (volleyballlover55):

A closed container has 4.02 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (1023 ⋅ 10–24) = 6.7134 5.69 grams, because (4.02 + 1.67) ⋅ (1023 ⋅ 10–24) = 5.69 0.67 grams, because (4.02 • 1.67) ⋅ (1023 ⋅ 10–24) = 6.7134 • 10–1 0.57 grams, because (4.02 + 1.67) ⋅ (1023 ⋅ 10–24) = 5.69 ⋅ 10–1

OpenStudy (volleyballlover55):

@Bobo-i-bo

OpenStudy (volleyballlover55):

@surjithayer

OpenStudy (anonymous):

\[(4.02*10^{23})*(1.67*10^{-24})=(4.02*1.67)*10^{23-24}=(4.02*1.67)*10^{-1}=?\]

OpenStudy (volleyballlover55):

um, 0.67134?

OpenStudy (anonymous):

this is equivalent to third option.

OpenStudy (volleyballlover55):

ok ! :)

OpenStudy (volleyballlover55):

What is the rational number equivalent to 3 point 24 with a bar over 24? 3 and 6 over 97 3 and 28 over 99 3 and 8 over 33 3 and 10 over 32

OpenStudy (anonymous):

you know geometric series.

OpenStudy (volleyballlover55):

no

OpenStudy (anonymous):

if not let